OpenStudy (anonymous):

for the following sequence {a^n}n>=1, find limn->infinity a^n or state why it does not exist:

OpenStudy (anonymous):

\[1. a _{n} = \log(n) - \log(n+1)\]

OpenStudy (anonymous):

2. \[a _{n} = \log(e+e ^{n}) \div 3n\]

OpenStudy (anonymous):

3. \[a _{n} = n!/n ^{n}\]

OpenStudy (anonymous):

4. \[a _{n} = (e ^{n}-e ^{-n})/(e ^{n}+e ^{-n})\]

OpenStudy (anonymous):

5. \[a _{n} = \sin^{-1} ((1-n ^{-1})^{3}) + \cos^{-1} ((1-n ^{-1})^{3})\]

OpenStudy (anonymous):

6. \[a _{n} = e ^{-n ^{2}} \int\limits_{0}^{n} e ^{t ^{2}}\]

OpenStudy (anonymous):

7. \[a _{n} = [\log(n)]^{\log(1-n ^{-1})}\]